# MambaTAD: When State-Space Models Meet Long-Range Temporal Action Detection

### v1

@@ B Error Analysis
In this section, we evaluate the detection results on THUMOS14 using the tool from [alwassel2018diagnosing]. The analysis focuses on three main aspects: sensitivity to varying lengths, False Positives (FP), and False Negatives (FN). Specifically, several characteristic metrics were defined in THUMOS14 in [alwassel2018diagnosing], including coverage, length, and the number of instances. Coverage is divided into five bins: Extra Small (XS: $(0,0.02]$), Small (S: $(0.02,0.04]$), Medium (M: $(0.04,0.06]$), Large (L:$(0.06,0.08]$), and Extra Large (XL: $(0.08,1.0]$), and length is organized into five groups: Extra Small (XS: $(0,3]$), Small (S: $(3,6]$), Medium (M: $(6,12]$), Long (L: $(12,18]$), and Extra Long (XL: >18>18). Besides, the number of instances represents the total count of occurrences of the same class within a video. This count is further categorized into four groups: Extra Small (XS: $1$), Small (S: $[2,40]$), Medium (M: $[40,80]$), and Large (L: >80>80).

### v2

@@ B Error Analysis
In this section, we evaluate the detection results on THUMOS14 using the tool from [2]. The analysis focuses on three main aspects: sensitivity to varying lengths, False Positives (FP), and False Negatives (FN). Specifically, several characteristic metrics were defined in THUMOS14 in [2], including coverage, length, and the number of instances. Coverage is divided into five bins: Extra Small (XS: $(0,0.02]$), Small (S: $(0.02,0.04]$), Medium (M: $(0.04,0.06]$), Large (L:$(0.06,0.08]$), and Extra Large (XL: $(0.08,1.0]$), and length is organized into five groups: Extra Small (XS: $(0,3]$), Small (S: $(3,6]$), Medium (M: $(6,12]$), Long (L: $(12,18]$), and Extra Long (XL: >18>18). Besides, the number of instances represents the total count of occurrences of the same class within a video. This count is further categorized into four groups: Extra Small (XS: $1$), Small (S: $[2,40]$), Medium (M: $[40,80]$), and Large (L: >80>80).